# LG-VQ: Language-Guided Codebook Learning

### v1

@@ 2.1 Vector Quantization for Image Generation
Vector quantization (VQ) is designed to learn a codebook, which aims to encode continuous image features into a discrete sequence. Then, the learned codebook can be utilized for various downstream tasks. Oord et al. van2017neural first propose a novel VQ method called VQ-VAE. This method innovatively replaces the prior distribution of Variational Autoencoder (VAE) with a discrete deterministic distribution (i.e., a codebook). To further improve the performance of VQ, various models are proposed to learn a more expressive codebook esser2021taming ; yu2021vitvq ; chang2022maskgit ; lee2022autoregressive ; Dong23PeCo ; huang2023towards ; huang2023not ; Lin23Catch . For example, VQ-GAN esser2021taming addresses the issue of image blur generated by VQ-VAE through the introduction of an adversarial training loss. However, the above methods do not tackle the codebook collapse issue. To address the issue, many novel methods are proposed from the perspective of regularization RazaviOV2019Generating , codebook update Zheng2023CVQ . Recently, inspired by the large language models (LLMs), instead of mapping images to the visual code tokens, some works attempt to map the images to the word tokens of LLMs by viewing images as “foreign languages” liu2023lqae ; yu2023spae ; zhu2024beyond . However, because of the inherent differences between vision and language, these works have difficulty assigning correct semantic words to images.
@@ 3.2.1 Semantic Alignment Module
$\mathcal{L}_{gsa}=-\sum_{i\in\mathcal{B}}\log\frac{\exp(s(e_{CLS}^{i},e_{EOT}^{i}))}{\sum_{j\in\mathcal{B}}\exp(s(e_{CLS}^{j},e_{EOT}^{j}))}.$ (5)
@@ 4.1 Experimental Settings
Dataset. We evaluate our method on four public datasets, including TextCaps sidorov2019textcaps , CelebA-HQ liu2015cele , CUB-200 wah2011cub , and MS-COCO lin2014coco .
Implementation Details. Following VQ-GAN esser2021taming , all images are reshaped 256 $\times$ 256 for reconstruction and generation. Down-sampling factor $f$ is set to 16. The codebook size $K$ is 1024. The batch size is 8. In our experiments, we maintain consistent parameter settings between our method LG-VQ and the chosen backbone networks (i.e., VQ-VAE van2017neural , VQ-GAN esser2021taming , and CVQ Zheng2023CVQ ) for a fair comparison. For each image, we randomly select a text from multi-text for training.
@@ 4.4.1 Image Generation
Following esser2021taming ; rombach2022high ; gu2022vqdiffusion , we conduct two image generation downstream tasks (i.e., text-to-image and semantic image synthesis) to fully validate the effectiveness of the learned codebook on CelebA-HQ.
Quantitative Evaluation. In Table 4.4.2, we compare our LG-VQ with the backbone methods on CelebA-HQ datasets for text-to-image and semantic image synthesis tasks. The results yield several key observations. Firstly, our LG-VQ method outperforms baseline methods in both tasks, affirming our method’s effectiveness. Secondly, the performance improvement of our method on the text-to-image task exceeds that on the semantic image synthesis task, which is reasonable due to the incorporation of pre-trained text knowledge enabling a comprehensive understanding of the text. This further confirms the advantage of introducing pre-trained text.
Qualitative Evaluation. We provide some synthesis examples comparing the results of our LG-VQ with baseline methods in Fig. 8, showing the performance in the text-to-image task. From the figure, we can see our method not only comprehensively understands the given text conditions but also excels in generating realistic images compared with baseline methods. For instance, our method can capture the “glasses”, “man”, “long black hair”, and “no beard” key attributions.
@@ 4.4.2 Visual Text Reasoning
Image captioning. Following mokady2021clipcap , we conduct the image captioning task on the CUB-200 dataset. We compare a recent work V2L Tokenizer zhu2024beyond accepted by CVPR2024, which maps images to the word tokens of LLMs. We select VQ-GAN as our backbone network. The results are shown in Table 7. For the results, we can see that our LG-VQ method outperforms the performance of VQ-GAN. This is reasonable because the pre-trained text provides rich context and relationship semantics for codebook learning, which verifies our motivation for learning a text-aligned codebook to improve the performance of the codebook on cross-modal tasks. On the other hand, the V2L Tokenizer cannot achieve very good performance because it is difficult to assign correct semantic language tokens to images. Compared with the V2L Tokenizer, our method utilizes pre-trained text semantics as supervised information. Its advantage is can make the codebook learn semantic information consistent with the text (i.e., learning a text-aligned codebook). And, our method is model-agnostic, which can be easily integrated into existing VQ models.
Visual Question Answering. We select VQ-GAN as the baseline. We conduct the VQA task on the COCO-QA ren2015cocoQA using the codebook trained on the MS-COCO dataset. The results are shown in Table 6. From the results, we can see that our LG-VQ method significantly improves the performance of VQ-GAN on VQA task (approximately 8.32$\uparrow$ on Accuracy). That is reasonable due to we introduce pre-trained text semantics to enable us to obtain a codebook aligned with the text, which is helpful for comprehensively understanding the given text question. This confirms our motivation and the effectiveness of our method.
@@
Setting VQ-VAE VQ-GAN CVQ
w/o LG w LG w/o LG w LG w/o LG w LG
text-to-image 49.51 49.36 15.29 12.61 13.23 12.33
image synthesis 48.72 48.23 11.53 11.46 11.04 11.03
@@
VQ-GAN 37.82 $\pm$ 0.97 82.06 $\pm$ 0.54
VQ-GAN+LG 40.97 $\pm$ 0.13 83.56 $\pm$ 0.11
Table 7: Results of image captioning on CUB-200 datasets.
@@
V2L Tokenizer (CVPR24) zhu2024beyond 1.59 30.65 25.76 104.14
VQ-GAN+LG 1.69 34.73 25.78 102.77
@@ A.1 Experiment Details
Semantic image synthesis: All models follow the default setting of VQ-GAN-Transformer [^1] https://github.com/CompVis/taming-transformers. Specifically, the vocabulary size, embedding number, and input sequence length are 1024, 1024, and 512, respectively. The layers and heads of the transformer are both 16. The training epoch is 30 for all models. The experiments are conducted on 1 4090 GPU with a batch size of 15 and one day of training time.
@@
A.2 More Examples and Qualitative Results
We provide more examples of image reconstruction in Fig. 10, image-to-text retrieval in Fig. 11 and Fig. 12. We also provide more image synthesis results in Fig. 13 for semantic image synthesis, and text-to-image synthesis in Fig. 14. We provide some examples of image captioning in Fig. 15 and VQA in Fig. 16. Moreover, we also apply our method of image completion on CelebA-HQ, generating images based on masked images. The results are shown in Fig. 17.

### v2

@@ 2.1 Vector Quantization for Image Generation
Vector quantization (VQ) is designed to learn a codebook, which aims to encode continuous image features into a discrete sequence. Then, the learned codebook can be utilized for various downstream tasks. Oord et al. van2017neural first propose a novel VQ method called VQ-VAE. This method innovatively replaces the prior distribution of Variational Autoencoder (VAE) with a discrete deterministic distribution (i.e., a codebook). To further improve the performance of VQ, various models are proposed to learn a more expressive codebook esser2021taming ; yu2021vitvq ; chang2022maskgit ; lee2022autoregressive ; Dong23PeCo ; huang2023towards ; huang2023not ; Lin23Catch . For example, VQ-GAN esser2021taming addresses the issue of image blur generated by VQ-VAE through the introduction of an adversarial training loss. However, the above methods do not tackle the codebook collapse issue. To address the issue, many novel methods are proposed from the perspective of regularization RazaviOV2019Generating , codebook update Zheng2023CVQ , codebook transfer zhang2024vqct . Recently, inspired by the large language models (LLMs), instead of mapping images to the visual code tokens, some works attempt to map the images to the word tokens of LLMs by viewing images as “foreign languages” liu2023lqae ; yu2023spae ; zhu2024beyond . However, because of the inherent differences between vision and language, these works have difficulty assigning correct semantic words to images.
@@ 3.2.1 Semantic Alignment Module
$\mathcal{L}_{gsa}=-\sum_{i\in\mathcal{B}}\log\frac{\exp(s(e_{CLS}^{i},e_{EOT}^{i}))}{\sum_{j\in\mathcal{B}}\exp(s(e_{CLS}^{i},e_{EOT}^{j}))}.$ (5)
@@ 4.1 Experimental Settings
Dataset. We evaluate our method on four public datasets, including TextCaps sidorov2019textcaps , CelebA-HQ liu2015cele , CUB-200 wah2011cub , and MS-COCO lin2014coco . For CelebA-HQ, CUB-200, and MS-COCO datasets, we use publicly available image captions, CelebA-HQ from xia2021tedigan , CUB-200 from reed2016learning , MS-COCO from chen2015microsoft
Implementation Details. Following VQ-GAN esser2021taming , all images are reshaped 256 $\times$ 256 for reconstruction and generation. Down-sampling factor $f$ is set to 16. The codebook size $K$ is 1024. The batch size is 8. In our experiments, we maintain consistent parameter settings between our method LG-VQ and the chosen backbone networks (i.e., VQ-VAE van2017neural , VQ-GAN esser2021taming , and CVQ Zheng2023CVQ ) for a fair comparison. For each image, we randomly select a text from multi-text for training. Since our method introduces additional text and pre-trained CLIP model, for a fair comparison, we select VQCT zhang2024vqct as the baseline for various downstream tasks. VQCT extracts many visual-related words from a large amount of text and designs a novel codebook transfer network based on the pre-trained CLIP model to learn the visual codebook.
@@
Table 2: Ablation study of our three loss functions on TextCaps and CUB-200.
Setting TextCaps CUB-200
[…]
(i) Baseline(VQ-GAN) 24.08 3.63
(ii) + $\mathcal{L}_{gsa}$ 23.01 3.39
(iii) + $\mathcal{L}_{mtp}$ 21.54 3.49
(iv) + $\mathcal{L}_{mtp}$ + $\mathcal{L}_{ras}$ 20.77 3.32
(v) + $\mathcal{L}_{mtp}$ + $\mathcal{L}_{gsa}$ 20.46 3.34
(vi) + $\mathcal{L}_{mtp}$ + $\mathcal{L}_{gsa}$ + $\mathcal{L}_{ras}$ 20.35 3.08
Table 3: Results (Recall@1) of masked word prediction on CelebA-HQ and CUB-200. “Mask-1” denotes that text is randomly masked one word.
Dataset Recall@1
CelebA-HQ Mask-1 99.55
Mask-3 99.24
CUB-200 Mask-1 83.65
Mask-3 80.17
@@ 4.4.1 Image Generation
Following esser2021taming ; rombach2022high ; gu2022vqdiffusion , we conduct image generation downstream tasks (i.e., text-to-image, semantic synthesis, unconditional generation, and image completion) to fully validate the effectiveness of the learned codebook on CelebA-HQ.
Text-to-Image. In Table 4.4.2, we compare our LG-VQ with the state-of-the-art models on CelebA-HQ dataset for text-to-image. From the results, our LG-VQ method outperforms baseline methods by a large margin. This is reasonable due to the incorporation of pre-trained text knowledge enabling a comprehensive understanding of the text, which suggests our method’s effectiveness. Moreover, we provide some synthesis examples comparing the results of our LG-VQ with baseline methods in Figure 8, showing the performance in the text-to-image task. From the figure, we can see our method not only comprehensively understands the given text conditions but also excels in generating realistic images compared with baseline methods. For instance, our method can capture the “glasses”, “man”, “long black hair”, and “no beard” key attributions.
Semantic Synthesis. Following esser2021taming , we compare with existing semantic synthesis models in Table 6. Our method achieves the best performance, which suggests our method’s effectiveness. We provide some examples in Appendix Figure 13.
Unconditional Generation and Image Completion. Following esser2021taming , we conduct unconditional image generation and Image Completion on CelebA-HQ dataset, as shown in Table 7 and Table 10. From the results, we can see that our method can significantly improve the performance of VQ-GAN, which is reasonable because pre-trained text can provide rich semantic knowledge for learning more robust codebook representation. This suggests the effectiveness of our method. We provide some examples in Appendix Figure 18 and Figure 17.
@@ 4.4.2 Visual Text Reasoning
Image captioning. Following mokady2021clipcap , we conduct the image captioning task on the CUB-200 dataset. We compare two recent work V2L Tokenizer zhu2024beyond and VQCT zhang2024vqct . We select VQ-GAN as our backbone network. The results are shown in Table 11. For the results, we can see that our LG-VQ method outperforms the performance of VQ-GAN. This is reasonable because the pre-trained text provides rich context and relationship semantics for codebook learning, which verifies our motivation for learning a text-aligned codebook to improve the performance of the codebook on cross-modal tasks. On the other hand, the V2L Tokenizer and VQCT cannot achieve very good performance because it is difficult to assign correct semantic language tokens to images. Compared with the V2L Tokenizer, our method utilizes pre-trained text semantics as supervised information. Its advantage is can make the codebook learn semantic information consistent with the text (i.e., learning a text-aligned codebook). And, our method is model-agnostic, which can be easily integrated into existing VQ models.
Visual Question Answering. We select VQ-GAN and VQCT zhang2024vqct as the baseline. We conduct the VQA task on the COCO-QA ren2015cocoQA using the codebook trained on the MS-COCO dataset. The results are shown in Table 9. From the results, we can see that our LG-VQ method significantly improves the performance of VQ-GAN on VQA task (approximately 8.32$\uparrow$ on Accuracy). That is reasonable due to we introduce pre-trained text semantics to enable us to obtain a codebook aligned with the text, which is helpful for comprehensively understanding the given text question. This confirms our motivation and the effectiveness of our method.
Table 5: Results of text-to-image on CelebA-HQ.
Model Text-to-Image
[…]
Unite and Conqu chen2020uniter 26.09
Corgi zhou2023corgi 19.74
LAFITE zhou2022lafite 12.54
VQ-GAN 15.29
CVQ 13.23
VQ-GAN+LG 12.61
CVQ+LG 12.33
[…]
Model Semantic Synthesis
[…]
Reg-VQ Zhang2023Regularized 15.34
VQCT zhang2024vqct 14.47
VQ-GAN 11.53
CVQ 11.04
VQ-GAN+LG 11.46
CVQ+LG 11.03
4.4.3 Visual Grounding
We conduct a visual grounding task on refcoco dataset yu2016refcoco to validate the effectiveness of the learned MS-COCO’s codebook. Following the same metric used in deng2021transvg , a prediction is right if the IoU between the grounding-truth box and the predicted bounding box is larger than 0.5. We select VQ-GAN and VQCT zhang2024vqct as the baseline. The results are shown in Table 8. From the results, we can see that the performance of our method consistently outperforms VQ-GAN and VQCT, which suggests its effectiveness. We also provide a qualitative comparison in Appendix Figure 19. For the experimental setting, please refer to Appendix A.1.
@@ 4.4.3 Visual Grounding
Model CelebA-HQ
[…]
Style ALAE parmar2021dual 19.2
DC-VAE pidhorskyi2020adversarial 15.8
VQ-GAN 10.2
LG-VQ 9.1
[…]
Model Visual Grounding
Accuracy(0.5$\uparrow$
VQ-GAN 9.14
VQCT zhang2024vqct 9.46
LG-VQ 9.62
@@
VQCT zhang2024vqct 40.42 82.06
VQ-GAN 37.82 83.22
LG-VQ 40.97 83.56
[…]
Model CelebA-HQ
[…]
VQ-GAN 9.02
LG-VQ 8.14
Improve 9.76%
Table 11: Results of image captioning on CUB-200.
@@
V2L Tokenizer zhu2024beyond 1.59 30.65 25.76 104.14
VQCT zhang2024vqct 1.38 26.50 24.63 98.22
LG-VQ 1.69 34.73 25.78 102.77
@@ 5 Conclusions
Acknowledgement
This work was supported by the Shenzhen Peacock Program under Grant No. ZX20230597, NSFC under Grant No. 62272130 and Grant No. 62376072, and the Shenzhen Science and Technology Program under Grant No. KCXFZ20211020163403005. It was also supported by the Major Key Project of PCL (PCL2023A08) and the National Science Foundation of China: Multi-source Cross-platform Video Analysis and Understanding for Intelligent Perception in Smart City: U20B2052.
@@ A.1 Experiment Details
Semantic image synthesis, unconditional generation and image completion: All models follow the default setting of VQ-GAN-Transformer [^1] https://github.com/CompVis/taming-transformers. Specifically, the vocabulary size, embedding number, and input sequence length are 1024, 1024, and 512, respectively. The layers and heads of the transformer are both 16. The semantic image synthesis experiments are conducted on 1 4090 GPU with a batch size of 15 and one day of training time. The unconditional generation and image completion experiments are conducted on 2 4090 GPUs with a batch size of 36 and one day of training time.
@@
Visual Grounding: The refcoco dataset [49] includes 19,994 images with 50,000 referred objects. Each object has more than one referring expression, and there are 142,210 referring expressions in this dataset. There are two commonly used split protocols for this dataset. One is RefCOCOg-google [25], and the other is RefCOCOgumd [28]. We follow RefCOCOgumd [28] to split the dataset. The train set has 42,404 expressions, the validation set has 3,811 expressions, and the test set has 3,785 expressions. Following [5], we concatenate the image codes and text tokens and feed them into a learnable transformer with coordinate regression layers (i.e., FNN) to predict the object box. The training epoch is 100 for all models. The experiments are conducted on 2 4090 GPUs with a batch size of 30 and several hours of training time.
@@
A.2 Experimental comparison with VQCT
We provide comparisons with VQCT [51] for image reconstruction in Table 12. Moreover, we also provide experimental results on further integrating our method into VQCT to verify its effectiveness and versatility. The results are shown in Table 13.
Table 12: Comparison of VQCT [51] and our method on reconstruction
Model Codebook Size #Tokens CelebA-HQ CUB-200 MS-COCO
VQCT 6207 512 5.02 2.13 9.82
VQ-GAN+LG 1024 256 5.34 3.08 10.72
CVQ+LG 1024 256 4.90 3.33 9.69
Image Reconstruction VQA
[…]
VQCT 9.82 40.42
VQCT+LG 9.57 40.64
Table 13: Comparison of reconstruction and VQA on VQCT and VQCT+LG on the MS-COCO dataset.
A.3 More Examples and Qualitative Results
We provide more examples of image reconstruction in Fig. 10, image-to-text retrieval in Fig. 11 and Fig. 12. We also provide more image synthesis results in Fig. 13 for semantic image synthesis, and text-to-image synthesis in Fig. 14. We provide some examples of image captioning in Fig. 15 and VQA in Fig. 16. We provide some examples of unconditional generation in Fig. 18, and image completion in Fig. 17. We also provide a qualitative comparison of visual grounding in Fig. 19.
@@ A.3 More Examples and Qualitative Results
Figure 18: Examples of unconditional image generation on CelebA-HQ based on VQ-GAN+LG.
Figure 19: Examples of visual grounding on refcoco. Blue boxes are the ground-truth, red boxes are the model predictions.